# Soft Reasoning on Uncertain Knowledge Graphs

### v1

@@ 3.2 Soft queries: formal definitions
Challenge of open world assumption. Similar to the study in first-order query answering, The answer utility for soft query $\Phi(y)$ in the open world $\mathcal{UKG}_{\rm OWA}=(\mathcal{KG},\mathcal{P})$ can be defined as:
$U_{\rm OWA}[\Phi,\mathcal{P}](s)=[\mathcal{P}\models_{s}\Phi(s/y)].$ (49)
We could see that the difference between $\mathcal{P}$ and $P$ causes the failure of applying classic methods on $P$ to estimate the utility function defined by $\mathcal{P}$. Similar to the necessity of ML generalization for logical queries on KG, the machine learning approach is also essential to mitigate the open world assumption in UKG.
Uncertain Knowledge Graph Embeddings (UKGE) were proposed as the extension of link predictors on traditional KG (Chen et al., 2019, 2021a). In general, UKGEs are trained in $\mathcal{UKG}$ but aim to approximate $\mathcal{UKG}_{\rm OWA}$. They are abstracted as the following interface with more introduction provided in Appendix E.
Definition 10 (UKGE).
A UKGE $(\mathcal{KG},\hat{P})$ parameterizes a differentiable function $\hat{P}:\mathcal{E}\times\mathcal{R}\times\mathcal{E}\mapsto[0,1]$, which approximates the function $\mathcal{P}$ for $\mathcal{UKG}_{\rm OWA}$. We assume the following uniform error-bound:
$\Pr\left(\max_{(s,r,o)}\|\hat{P}(s,r,o)-\mathcal{P}(s,r,o)\|>\delta\right)<\varepsilon(\delta),$ (50)
where $\varepsilon(\delta)$ is the tail probability of the uniform error $\delta$.
The error bound $\varepsilon(\delta)$ is assumed to facilitate the error analysis of our algorithm. Theories to estimate $\varepsilon(\delta)$ and methods to eliminate $\varepsilon(\delta)$ are beyond the scope of this paper.
Notation remark. We can see that the confidence functions ($P$, $\mathcal{P}$, and $\hat{P}$) are the only different element to $\mathcal{UKG}$, $\mathcal{UKG}_{\rm OWA}$, and the UKGE. To avoid the abuse of notation, we use $P,\mathcal{P},\hat{P}$ to refer to the $\mathcal{UKG}$, $\mathcal{UKG}_{\rm OWA}$, and UKGE, respectively. Furthermore, we use $U$, $\hat{U}$. and $\mathcal{U}$ to represent corresponding utility function under the confidence function $P$, $\hat{P}$, and $\mathcal{P}$, respectively. For example, the $U[\Phi,P](s)$ can be denoted as $U[\Phi](s)$ for any soft query $\Phi$ and entity $s$.
We tell the uniqueness of SQUK setting from the differences in UKG and KG, soft queries against logical queries, and the challenges from two versions of open world assumptions. The difference is also demonstrated in Table 1.
3.3 Explaining soft queries: a case study of Example 1
We are now ready to fully elaborate Example 1 and Figure 1. Let the Has be the relation describing a candidate who has a skill. The soft queries for candidates for Junior Software Developers (JSD) and Principle Investigators (PI) in ML are written in our notation
$\phi_{\rm JSD}(y)=$$\lnot(\textsc{Has}(y,\textsc{Lead}),(0.7,1))\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}(\textsc{Has}(y,\textsc{Dev}),(0.5,3)),$
@@ 3.3 Explaining soft queries: a case study of Example 1
$\phi_{\rm PI}(y)=$ $(\textsc{Has}(y,\textsc{Lead}),(0.7,3))\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}(\textsc{Has}(y,\textsc{ML}),(0.9,1)).$
For the JSD job, Person 1 is overqualified because his leadership style suits more senior ones, while Person 2 is the perfect candidate. For the PI job, Person 1 and 2 are not suitable because of their limited ML research skills. Person 3 is suitable for ML research, but the closed world knowledge (solid lines in Figure 1) is not sufficient to check whether s/he is a good developer. With the machine learning model, the tendency for candidates not to possess Dev and ML skills simultaneously can be learned. Therefore, the ML model might estimate Person 3 has dev skill with the confidence value 0.1 (indicated in the dashed line in Figure 1), making him/her not a good candidate for the JSD job.
@@
0:  Input soft query graph ${G}_{\phi}$ and initialize $\{C_{z}\}$.
0:  Output utility vector $\hat{U}(G_{\phi},\{C_{z}\})$
@@
4 Methodology
In this section, we propose Soft Reasoning with calibrated Confidence values (SRC) to enable reasoning with different query structures and soft requirements. As an inference method based on confidence values, SRC  is differentiable and parallelized as its predecessors (Bai et al., 2023; Yin et al., 2024) and can seamlessly work with UKGE, which mitigates the open world assumption. Under a mild assumption of UKGE error bound in Equation (50), our error analysis suggests that the inference error of SRC  is controllable with the growing complexity of query structure. To further reduce the error, we propose two orthogonal calibration strategies: Debiasing (D) and Learning (L).
4.1 Forward inference
The goal of the inference in SRC  is to solve the evaluation problem $[\mathcal{P}\models_{s}\Phi(s/y)]$ for all $s\in\mathcal{E}$, where $\mathcal{P}$, as mentioned before, is the confidence function with OWA, $\models_{s}$ is the soft entailment operator, and $\Phi(s/y)$ is the soft query with free variable $y$ substituted by $s$, as defined in Section 3.
Remark 2.
For simplicity, we only discuss Soft Conjunctive Queries (SCQs) $\phi$ in Equation (42) because the soft disjunction can be addressed by the first soft entailment rule after solving all SCQs independently.
To demonstrate our algorithm, we introduce SCQ graphs.
Definition 11 (SCQ graph).
Given an SCQ $\phi=\exists x_{1},...,x_{n}.\psi_{1}\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}\cdots\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}\psi_{m}$, the SCQ graph $G_{\phi}$ is defined by tuples induced by atomic soft formulas $\psi_{1},...,\psi_{m}$: $G_{\phi}=\{(h_{i},r_{i},t_{i},\alpha_{i},\beta_{i},\textsc{Neg}_{i})\}_{i=1}^{m}$, where $(h_{i},r_{i},t_{i},\alpha_{i},\beta_{i})$ is naturally induced by expanding $\psi_{i}=(r_{i}(h_{i},t_{i}),(\alpha_{i},\beta_{i}))$ or $\psi_{i}=(\lnot r_{i}(h_{i},t_{i}),(\alpha_{i},\beta_{i}))$. $\textsc{Neg}_{i}=\textsc{True}$ if $\psi_{i}$ is negated and $\textsc{Neg}_{i}=\textsc{False}$ otherwise.
If $h$ (or $t$) is a variable, we say the corresponding node in $G_{\phi}$ is a variable node. $V(G_{\phi})$ indicates the set of all variable nodes in $G_{\phi}$ If $h$ (or $t$) is an entity, we say the corresponding node is an constant node. The leaf node is a node that is only connected to one other node in the query graph. The main paper discusses soft queries in which the query graphs are acyclic simple graphs. Cases with the complete case (cycles and self-loops) are detailed in Appendix D.
Recall that the utility vector $\mathcal{U}$ represents the utility of the substitution $[\mathcal{P}\models_{s}\phi(s/y)]$. With the estimated uncertain KG $\hat{P}$ to approximate UKG under OWA, we can calculate the estimated utility of $\hat{U}(s)=[\hat{P}\models_{s}\phi(s/y)]$.
SRC  derives the utility vector $\hat{U}$ with efficient inference. The core idea is to progressively prune the edges of the soft query graph while preserving the final utility vector unchanged. The state vectors are employed to track confidence values during the inference. Specifically, each variable node $z\in G_{\phi}$ is described by a state vector $C_{z}\in\mathbb{R}^{|\mathcal{E}|}_{+}$. $(G_{\phi},\{C_{z}:z\in V(G_{\phi})\})$ denotes an SCQ graph with intermediate results of inference.
We define equivalent transformations $T$:
$T(G_{\phi},\{C_{z}:z\in V(G_{\phi})\})=(G_{\psi},\{C_{z}^{\prime}\}:z\in V(G_{\psi})),$
where $G_{\psi}$ is a subgraph of $G_{\phi}$ (with at least one edge eliminated), $C_{z}^{\prime}$ is the updated state vector, and $T$ guarantees the final utility vecotr $\hat{U}$ unchanged.
Two lemmas are presented to induce two equivalent transformations, denoted as $T_{e}$ and $T_{l}$, respectively. The proof can refer to Appendix D.
Lemma 1.
For each constant node in an SCQ graph $G_{\phi}$, an $O(|\mathcal{E}|)$ transformation $T_{c}$ exists to remove it.
Lemma 2.
For each leaf node in an SCQ graph $G_{\phi}$, an $O(|\mathcal{E}|^{2})$ transformation $T_{l}$ exists to remove it.
@@ 4.1 Forward inference
The procedure of SRC for simple and acyclic SCQ graphs is presented in Algorithm 1. We first remove the constant nodes by Lemma 1. Then, we can find the leaf node and remove it step by step according to Lemma 2. Until the soft query only contains a free variable, the state vector $C_{y}$ of the free variable is the desired utility vector.
Figure 2: Query structures of selected query types. The white, yellow, and red circles represent constant, existential, and free nodes, respectively. Negative constraints are represented by red edges, while positive constraints are represented by black edges. Following the naming convention in previous work (Ren & Leskovec, 2020; Yin et al., 2024), we use ”p” for projection, “i” for intersection, “n” for negation, “m” for multi-edge, and “l” for existential leaf.
Complexity analysis. The space complexity of inference algorithm is $O(|\mathcal{R}||\mathcal{E}|^{2})$. Additionally, let $n_{e}$ represent the number of edges involving the existential variable and other variables, and $n_{r}$ denote the count of remaining edges. When solving queries without cycles, the time complexity is $O(n_{e}|\mathcal{E}|^{2}+n_{r}|\mathcal{E}|)$. The result of complexity is the same with previous works (Bai et al., 2023; Yin et al., 2024).
@@ 4.2 Error analysis
We first look into the error of each soft atomic query:
@@
For a soft atomic query $\psi=(a,(\alpha,\beta))$, let the uniform inference error be
$\max_{\psi,s\in\mathcal{E}}\|\hat{U}[\psi](s)-\mathcal{U}[\psi](s)\|=\epsilon(\alpha,\beta)$
@@
$\Pr\left(\epsilon(\alpha,\beta)>\delta\right)$
$<\varepsilon(\frac{\delta}{\beta})+(1-\varepsilon(\frac{\delta}{\beta}))\int_{0}^{1}\varepsilon(|\alpha-\xi|)f(\xi){\rm d}\xi.$ (51)
@@
For a soft conjunctive query $\phi=\exists x_{1},...,x_{n}.\psi_{1}\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}\cdots\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}\psi_{m}$, where $\psi_{i}=(a_{i},(\alpha_{i},\beta_{i}))$, and any entity $s\in\mathcal{E}$, the error accumulated is at most linear:
$\|\hat{U}[\phi](s)-\mathcal{U}[\phi](s)\|\leq\Sigma_{i=1}^{m}\epsilon(\alpha_{i},\beta_{i}).$ (52)
This conclusion ensures that there is no catastrophic cascading error in our forward inference algorithm.
The proof of all the above theorems can refer to Appendix J.

### v2

@@ 3.1 Syntax and semantic
  1. 1.
    If $\phi$ is the substituted soft atomic query $(s,r,o,\alpha,\beta)$, then $U(\phi,P)=\beta[P(s,r,o)]_{\alpha}$;
  2. 2.
    If $\phi$ is the negation of the substituted soft atomic $\lnot(s,r,o,\alpha,\beta)$, then $U(\phi,P)=\beta[1-P(s,r,o)]_{\alpha}$;
  3. 3.
    If $\phi=\exists x_{i}\psi(y;x_{i})$ is the soft query involving existentially quantified variables, then $U(\phi,P)=\oplus_{s\in\mathcal{E}}U(\phi(y;s/x_{i}),P)$;
  4. 4.
@@
    If $\phi$ is the conjunctive query $(\phi_{1}\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}\phi_{2})$, then $U(\phi,P)=U(\phi_{1},P)\otimes U(\phi_{2},P)$.
  5. 5.
    If $\Phi$ is the disjunctive query $(\Phi_{1}\mathchoice{\mathbin{\ooalign{$\displaystyle\lor$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\lor$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\lor$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\lor$\crcr$\scriptscriptstyle\bigcirc$}}}\Phi_{2})$, then $U(\Phi,P)=U(\Phi_{1},P)\oplus U(\Phi_{2},P)$.
[…]
Definition 6 (Utility vector).
Given a confidence function $P$ induced by an uncertain knowledge base and a soft query $\phi$, the utility vector of soft query $\Phi$, denoted as $\mathbf{u}\in\mathbb{R}^{|\mathcal{E}|}$, is defined as:
$\mathbf{u}_{i}=U(\Phi(s_{i}/y),P),$ (39)
where $s_{i}$ denotes the entity indexed by $i$.
3.2 Example to explain the soft queries, as well as the necessity and importance
Soft queries are a powerful tool for modeling candidate searches across various job positions, enabling nuanced assessments of qualifications. To illustrate this, we present an example of using soft queries to model candidate searches for two roles: Junior Software Developer (JSD) and Principal Investigator (PI) in machine learning. Let Has denote the relation describing a candidate’s possession of a skill, while LEAD, DEV, and ML represent leadership, development, and machine learning skills, respectively. Regarding leadership, the Principal Investigator places a significantly high emphasis on it, while the Junior Software Developer does not. This distinction is captured by the importance parameter $\beta$, which assigns greater weight to leadership in the query for the Principal Investigator. Additionally, the necessity parameter $\alpha$ serves as a threshold to filter out candidates who lack the required skills. The two roles can be modeled using the following soft queries, as further explained in Figure 1:
$\phi_{\rm JSD}(y)=$$\lnot(y,\textsc{Has},\textsc{Lead},0.7,1))\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}(y,\textsc{Has},\textsc{Dev},0.5,3),$
$\phi_{\rm PI}(y)=$ $(y,\textsc{Has},\textsc{Lead},0.7,3)\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}(y,\textsc{Has},\textsc{ML},0.9,1)).$
Beyond recruitment, soft queries provide an effective and tailored strategy for modeling logical queries on uncertain KGs. This capability can extend the application of logical queries to uncertain KGs, enabling use cases such as product recommendation (Bai et al., 2024b) and understanding user intentions (Bai et al., 2024a).
3.3 Soft query graph and utility vector
Definition 7 (Soft query graph).
Given a soft query $\phi(y;x_{1},...,x_{n})=\exists x_{1},...,x_{n}.a_{1}\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}\cdots\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}a_{m}$, the soft query graph $G_{\phi}$ is defined by tuples induced by soft atomic formulas or its negation: $G_{\phi}=\{(h_{i},r_{i},t_{i},\alpha_{i},\beta_{i},\textsc{Neg}_{i})\}_{i=1}^{m}$, where $(h_{i},r_{i},t_{i},\alpha_{i},\beta_{i},\textsc{Neg}_{i}))$ is induced by $(h_{i},r_{i},t_{i},\alpha_{i},\beta_{i})$ or $\lnot(h_{i},r_{i},t_{i},\alpha_{i},\beta_{i})$. $\textsc{Neg}_{i}$ is the bool variable indicating if $a_{i}$ is negated.
If $h$ (or $t$) is a variable, we say the corresponding node in $G_{\phi}$ is a variable node. $V(G_{\phi})$ indicates the set of all variable nodes in $G_{\phi}$. If $h$ (or $t$) is an entity, we say the corresponding node is an constant node. The leaf node is a node that is only connected to one other node in the query graph. Compared to the operation tree in complex logical queries (Ren and Leskovec, 2020), the soft query graph can model any conjunctive soft queries.
4 Methodology
In this section, we propose Soft Reasoning with calibrated Confidence values (SRC) to facilitate reasoning with various query structures and soft requirements. SRC is a symbolic reasoning method that utilizes UKGE to provide confidence values. Since UKGE inevitably has prediction errors, we present a mild assumption regarding the UKGE error bound in Equation (41). Our error analysis over SRC indicates that the inference error is manageable as the complexity of the query structure increases. To further reduce this error, we introduce two orthogonal calibration strategies: Debiasing (D) and Learning (L).
4.1 Forward inference
Figure 2: A toy model illustrating the process of the SRC. Each variable node is assigned a state vector, which is updated through the algorithm as edges are removed. The final state vector of the free variable is the desired.
The main paper discusses soft queries in which the query graphs are acyclic simple graphs. Cases with the complete case (cycles and self-loops) are detailed in Appendix H.
Given the soft query $\phi$, SRC  efficiently derives the utility vector $U(\phi,\hat{P})$ based on the confidence function $\hat{P}$ approximated by UKGE. The core idea is to progressively prune the edges of the soft query graph while preserving the constraints of the remaining edges, ensuring that the final utility vector remains unchanged. State vectors are used to record the constraints of the pruned edges during the inference process. Specifically, each variable node $z\in G_{\phi}$ is described by a state vector $C_{z}\in\mathbb{R}^{|\mathcal{E}|}$. The notation $(G_{\phi},\{C_{z}:z\in V(G_{\phi})\})$ denotes a soft query graph with state vectors.
We define equivalent transformations as $T$:
$T(G_{\phi},\{C_{z}:z\in V(G_{\phi})\})=(G_{\psi},\{C_{z}^{\prime}:z\in V(G_{\psi})\}),$ (40)
where $G_{\psi}$ is a subgraph of $G_{\phi}$ (with at least one edge eliminated), $C_{z}^{\prime}$ is the updated state vector, and $T$ guarantees the utility vector $\hat{\mathbf{u}}$ unchanged.
Two lemmas are presented to induce two equivalent transformations, denoted as $T_{e}$ and $T_{l}$, respectively. The proof can refer to Appendix H.
Lemma 1.
For each constant node in $G_{\phi}$, an $O(|\mathcal{E}|)$ transformation $T_{c}$ exists to remove it.
Lemma 2.
For each leaf node in $G_{\phi}$, an $O(|\mathcal{E}|^{2})$ transformation $T_{l}$ exists to remove it.
@@ 4.1 Forward inference
0:  Input soft query graph ${G}_{\phi}$ and initialize the state vectors $\{C_{z}\}$.
0:  Output utility vector $\hat{\mathbf{u}}(G_{\phi},\{C_{z}\})$.
@@
The above two transformations are constructive and can be applied to remove edges once the corresponding nodes are found. Until the soft query graph only contains a free variable node, the state vector of the free variable $C_{y}$ is the desired utility vector. The procedure of SRC for acyclic soft query graphs is presented in Algorithm 1 and one toy example of the execution is visualized in Figure 3. We first remove constant nodes, as these are commonly found and can be easily eliminated using Lemma 1. Subsequently, we identify and remove leaf nodes step by step, according to the Lemma 2, noting that the leaf node always exists for acyclic queries.
Complexity analysis. We begin with a rough estimation of the complexity. The space complexity of the inference algorithm is $O(|\mathcal{R}||\mathcal{E}|^{2})$. Let $n_{e}$ denote the number of edges involving existential variables and $n_{r}$ represent the remaining edges. For acyclic queries, the time complexity is $O(n_{e}|\mathcal{E}|^{2}+n_{r}|\mathcal{E}|)$. We further reduce complexity by leveraging sparsity, a natural characteristic of knowledge graphs. For space complexity, only non-zero values or those exceeding a threshold $\delta_{1}$ are stored and the average sparsity ratio of our constructed neural matrices is approximately $3.4\%$, reducing storage requirements by $97\%$. For time complexity, sparsity can be used to accelerate computation. Removing a leaf node $x$ typically involves $O(E^{2})$ operations, but by considering only rows where $C_{x}$ is non-zero, this reduces to $O(\mathcal{E}\cdot(|C_{x}>\delta_{2}|))$, where $|C_{x}>\delta_{2}|$ represents the non-zero number of $C_{x}$. Thus, the time complexity becomes $O(n_{e}\mathcal{E}\cdot\max_{x}(|C_{x}>\delta_{2}|))$, where $n_{e}$ is the number of edges involving existential variables.
@@ 4.2 Error analysis
The current model still exhibits significant prediction errors, with the mean absolute error on CN15K generally reaching around 0.2 (Chen et al., 2021a, b). To facilitate the error analysis of our proposed search algorithm, we introduce the error bound $\varepsilon(\delta)$ as the following:
Definition 8.
Let $\varepsilon$ be a function that maps the error $\delta$ to a tail probability which can uniformly bounds the error of $\hat{P}$ for some kind of norm:
$\Pr\left(\max_{(s,r,o)}\|\hat{P}(s,r,o)-\mathcal{P}(s,r,o)\|>\delta\right)<\varepsilon(\delta).$ (41)
We note that $\varepsilon(\delta)<1$ is guaranteed for all $\delta$ and $\varepsilon(\delta)$ decreases monotonically with $\delta$, though a better link predictor provides a tighter $\varepsilon(\delta)$. This definition provides a practical and flexible approach for analyzing the errors over uncertain KGs. We look into the error of each soft atomic query:
@@
For any soft atomic query $\psi=(h,r,y,\alpha,\beta)$, let the uniform inference error be
$\max_{\psi,s\in\mathcal{E}}\|U(\psi(s/y),\hat{P})-U(\psi(s/y),\mathcal{P})\|=\epsilon(\alpha,\beta)$
@@
$\Pr\left(\epsilon(\alpha,\beta)>\delta\right)<\varepsilon(\frac{\delta}{\beta})+(1-\varepsilon(\frac{\delta}{\beta}))\int_{0}^{1}\varepsilon(|\alpha-\xi|)f(\xi){\rm d}\xi.$
@@
For a soft conjunctive query $\phi=\exists x_{1},...,x_{n}.a_{1}\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}\cdots\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}a_{m}$, where $a_{i}=(h_{i},r_{i},t_{i},\alpha_{i},\beta_{i})$, and any entity $s\in\mathcal{E}$, the error accumulated is at most linear:
$\|U(\phi(s),\hat{P})-U(\phi(s),\mathcal{P})\|\leq\Sigma_{i=1}^{m}\epsilon(\alpha_{i},\beta_{i}).$ (42)
This conclusion ensures that there is no catastrophic cascading error in our forward inference algorithm. The proof of all the above theorems can refer to Appendix N.